OpenStudy (anonymous):

hey guys need a little help!!! (1/4,-1/2) and (3/4,5) this is asking me to find an equation of the line containing the given pair of points. please help!!

OpenStudy (amistre64):

subtract one form the other; slope it; and use the slope point form of an equation to define it

OpenStudy (anonymous):

first find the slope:\[m=\frac{y_2-y_1}{x_2-x_1}\] \[=\frac{-\frac{1}{2}-5}{\frac{3}{4}-\frac{1}{4}}\]

OpenStudy (anonymous):

hello amistre!

OpenStudy (amistre64):

(3/4 ,5) -(1/4,-1/2) ------------ 1/2, 11/2 ; slope = y/x = 11

OpenStudy (amistre64):

hi sat :)

OpenStudy (amistre64):

y = 11(x-1/4)-1/2

OpenStudy (anonymous):

\[=\frac{-\frac{11}{2}}{\frac{1}{2}}=-11\]

OpenStudy (anonymous):

i got -11 yes?

OpenStudy (amistre64):

y = 11x -11/4-1/2

OpenStudy (amistre64):

nope, both x and y are the same sign

OpenStudy (anonymous):

oops no sorry amistre has it

OpenStudy (anonymous):

you are right it is 11, my fault

OpenStudy (amistre64):

y = 11x -13/4 i thnk

OpenStudy (amistre64):

-44x +4y = -13 44x -4y = 13 maybe?

OpenStudy (amistre64):

the 'equation of a line' can take many forms; general, standard, slope int, slope point, etc..